What is the y-intercept form of the equation of the line that passes through points (2,3) and (-1,4)?
Answer by checkley71(8403)

SLOPE IS (4-3)/(-1-2) OR 1/-3 OR -1/3 THUS THE EQUATION IS
Y=-1/3X+b THEN 3=(-1/3)*2+b or 3=-2/3+b or b=3 2/3 or Y=-1/3X+3 2/3